# Received August 1, 2010; revised October 1, 2010; accepted for publication November 1, 2010

### v1

@@ Abstract
Uncertainty around predictions from a model due to the finite size of the development sample has traditionally been approached using classical inferential techniques. The finite size of the sample will result in the discrepancy between the final model and the correct model that maps covariates to predicted risks. From a decision-theoretic perspective, this discrepancy might affect the subsequent treatment decisions, and thus is associated with utility loss. From this perspective, procuring more development data is associated in expected gain in the utility of using the model. In this work, we define the Expected Value of Sample Information (EVSI) as the expected gain in clinical utility, defined in net benefit (NB) terms in net true positive units, by procuring a further development sample of a given size. We propose a bootstrap-based algorithm for EVSI computations, and show its feasibility and face validity in a case study. Decision-theoretic metrics can complement classical inferential methods when designing studies that are aimed at developing risk prediction models. predictive analytics; precision medicine; decision theory; value of information; bayesian statistics
@@ 1 Introduction
Risk prediction models that utilize patient characteristics to predict the risk of clinically relevant events are important enablers of precision medicine and individualized care (Kattan and others, 2016; Steyerberg, 2019; Van Calster and others, 2013). Typically, such models are classical regression-based or machine learning algorithms developed using a finite ’development’ sample from a relevant target population. Examples of such models are the Framingham Risk Score for predicting the risk of cardiovascular outcomes (D’Agostino Sr and others, 2008), or the Adnex score for predicting the risk of malignancy in ovarian tumors (Timmerman and others, 2016).
@@
Predictions from such models are inherently subject to uncertainty, primarily due to the finite size of the development sample (Harrell, 2015; Steyerberg, 2019). Conventionally, such uncertainty is communicated through standard frequentist methods around performance metrics of the model (e.g., 95% confidence interval around c-statistic or calibration line [DeLong and others, 1988]). Recent developments have focused on studying the stability of predicted risks, with proposals on the graphical and statistical presentations of variation in predicted risks (Riley and Collins, 2023). Similarly, in designing risk prediction development studies, contemporary approaches for sample size calculations are based on statistical inference on classical metrics of model performance. For example, Riley and others (2024) have proposed multi-criteria sample size rules targeting prediction error, calibration, and discrimination. These significant developments have provided much needed objectivity in the design of risk prediction development and validation studies.
@@
In addition to classical metrics of model performance around calibration and discrimination, the clinical utility of a model, expressed in terms of net benefit (NB), has become a standard metric of model performance (Vickers and Elkin 2006a; Efthimiou and others 2024). Among conventional metrics of model performance, net benefit (NB) is a decision-theoretic one, in the sense that the use of a risk prediction model is clinically justified if its expected NB is above the expected NB of the alternative decisions. Decision theory provides a fundamentally different approach towards uncertainty quantification than methods based on conventional statistical inference (Claxton, 1999). From the perspective of this theory, uncertainty results in the discrepancy between the predicted and correct (calibrated) risks, which in turn might result in an incorrect medical decision (a decision based on predicted risk that is different from when knowing the correct risk). Thus, from this perspective, uncertainty can be associated with a loss of clinical utility (Steyerberg, 2019). Such decision-theoretic approach towards evaluating the consequences of uncertainty in decisions, generally referred to as Value of Information (VoI) analysis, has been employed in various fields, such as decision-analytic modeling (Jackson and others, 2022; Jackson and others, 2019), environmental risk analysis (Yokota and Thompson, 2004; Weinstein, 1980; O’Hagan and others, 2006), and economic trials (Everest and others, 2023). Recently, such methods have been adapted to predictive analytics (Jackson and others, 2022; Vickers and others, 2023). Specifically, the expected value of perfect information (EVPI) has been applied to both development and validation phases of risk prediction models (Sadatsafavi and others, 2022; Sadatsafavi and others, 2023a). In the context of model development, EVPI quantifies the expected utility gain over a model developed using a finite sample by completely knowing the correct (strongly calibrated) model.
@@
However, uncertainty can never be entirely eliminated. Procuring more samples for model development is expected to reduce, but not eliminate, prediction uncertainty (Greenland, 2000). As such, the EVPI puts an upper limit on the expected gain from future validation studies. In decision theory, the expected value of sample information (EVSI) quantifies the potential value of collecting additional data to inform a specific research design (Claxton, 1999; Willan and Briggs, 2006). In health policy-making, EVSI has been applied to quantify the expected utility (net benefit) gain from conducting a clinical trial of a given design and sample size (Ades and others, 2004a). Such expected gain can be juxtaposed with the costs of the trial to optimize the sample size (Willan and others, 2012).
@@
The EVSI has recently been proposed for the validation phase of a model (evaluating the performance of a pre-specified model in a new population without changing its structure [Sadatsafavi and others, 2023b]). To the best of our knowledge, it has not yet been applied to the development phase of such models . Recent best practice standards recommend for not splitting the available data into development and validation, and instead use all the data for development [Collins and others, 2024]. As well, in many instances where a model is being validated in a new population, some form of model updating will be required (ranging from changing the intercept, to refitting the model in the new population [Binuya and others, 2022]). These mean in many practical instances, the investigators need to assess the NB of their model in the same data that is used to fit or update the model. As such, informing sample size calculations for model development around NB is a relevant pursuit.
@@ 2 Context and Notations
We focus on the development phase of a risk prediction model for a binary outcome (e.g., whether an individual with a heart attack will die in the next 30 days). We assume the investigator has had access to $n$ iid observations of predictors and response values from the relevant target population to develop a prediction model for the conditional risk of the outcome given a set of predictor values. Due to the finite size of the development sample, predictions from such a model might not be accurate even if the model form is correctly specified, resulting in potential loss of clinical utility. Procuring a further development sample can improve the accuracy of predictions, and reduce such a loss. The figure of merit is the expected gain in clinical utility from procuring a future development sample $D^{*}$ of size $n^{*}$ from the same target population, i.e., development EVSI. We assume the future sample will be exchangeable with the current sample (thus the investigator will naturally combine the current and future samples to refit their model). We quantify clinical utility in terms of net benefit (NB), as quantified in decision curve analysis (Vickers and Elkin, 2006b).
@@ 2.1 Notations
Let $D=(\mathbf{X},Y)=\{(\mathbf{X_{i}},Y_{i})\}_{i=1}^{n}$ be the current development sample where $\mathbf{X}$ represents the vector of predictors and $Y$ the binary outcome, and consequently $d=(\mathbf{x},y)=\{(\mathbf{x_{i}},y_{i})\}_{i=1}^{n}$ be its realization after observing the sample. The development data are presumed to obey the following joint distribution
$P_{\theta,\eta}(X,Y)=P_{\eta}(X)P_{\theta}(Y|X),$ (1)
where $\eta$ and $\theta$ are the parameters indexing, respectively, the marginal distribution of $X$ (i.e., the case-mix) and the conditional distribution of Y given X ($Y|X$, i.e., the response model). Given the binary nature of $Y$, let $\pi(x;\theta)=P(Y=1|X=x)$ be the risk of outcome for patient with characteristics $x$, and consequently, $\pi(X;\theta)=P(Y=1|X)$ be a random draw from the distribution of predicted risks specified by a model parameterized by $\theta$. Related, $\hat{\eta}$ and $\hat{\theta}$ estimate the case mix and model coefficients from the available data $d$, respectively. In what follows, we parametrically model $Y$, estimating $\theta$ (regression coefficients), in a general regression framework, while we adopt a non-parametric approach towards modeling case-mix, where $\hat{\eta}$ is the empirical distribution of $\mathbf{X}$.
@@ 2.2 Net benefit
We determine clinical utility by using NB, a widely recognized method for assessing the performance of risk prediction models due to its solid theoretical foundation and straightforward application Vickers and Elkin, 2006a. When converting continuous risk predictions into binary decisions, such as deciding whether or not to treat a patient, it is necessary to set a treatment threshold that aligns with the specific clinical context. This threshold should ideally reflect the relative importance of avoiding unnecessary treatments (false positives) versus ensuring necessary treatments (true positives). As explained by Vickers and Elkin, 2006a, this threshold effectively functions as a balancing point between true positives and false positives, allowing NB to be calculated by comparing the benefits of correct treatment decisions against the drawbacks of incorrect ones.
To compute NB, a particular risk threshold, denoted by $z$, is chosen. Patients with a predicted risk below $z$ are considered low-risk and are not treated, while those with a risk above $z$ are classified as high-risk and are treated. For those whose predicted risk is exactly $z$, there is no strong preference for treatment or non-treatment, as treating these patients would lead to a certain proportion being appropriately treated (true positives) and the rest being treated unnecessarily (false positives). This balance suggests that the negative impact of a false positive is weighted by $\frac{z}{1-z}$ in relation to a true positive. By using this threshold, the model’s true and false positive rates can be calculated, and NB can be expressed as the net gain in true positives after accounting for the cost of false positives. This approach provides a flexible way to evaluate NB across different thresholds $z$ and can be expressed as follows (dependency on $z$ omitted for simplicity)
NB of $\text{NB of model}=P(\text{True Positive})-P(\text{False Positive})\frac{z}{1-z}$ (2)
where, as mentioned earlier, $\frac{z}{1-z}$ represents the relative weight of a false-positive versus a true-positive classification and thus captures the harm-benefit trade-off at this threshold. In practice, the NB is often calculated for a plausible range of thresholds.
To make a decision whether the model is useful in this population, we shall compare its NB against the NB of at least two default decisions: treating all and treating none. The latter has 0 NB by definition, and the NB of treating all can be written as
@@
We index these three strategies as 0=treat none, 1=use model, and 2=treat all. Then the true NB function for a model indexed by $\hat{\theta}$ can generally be defined as (for brevity we drop the notation that indicates the NB quantities are dependent on $z$):
$NB\left(i;\hat{\theta},\eta,\theta\right)=\left\{\begin{array}[]{ll}0&i=0\text{ (treat none)}\\ \mathbb{E}\left\{I\left(\pi\left(X,\hat{\theta}\right)\geq z\right)\left[\pi(X,\theta)-(1-\pi(X,\theta))\frac{z}{1-z}\right]\right\}&i=1\text{ (use model)}\\ \mathbb{E}\{\pi(X,\theta)-(1-\pi(X,\theta))\frac{z}{1-z}\}&i=2\text{ (treat all)}\end{array}\right.,$ (4)
@@
In the current development sample $d$ of size $n$, given $\theta$ and $\hat{\theta}$ while plugging in the empirical estimate of $\eta$ (a weight of $1/n$ assigned to each observation), a consistent estimate of NBs can be written as:
$\hat{NB}\left(i;\hat{\theta},\theta\right)=\left\{\begin{array}[]{ll}0&i=0\\ \frac{1}{n}\sum_{j=1}^{n}\left(I\left(\pi\left(x_{j},\hat{\theta}\right)>z\right)\left[\pi(x_{j},\theta)-(1-\pi(x_{j},\theta))\frac{z}{1-z}\right]\right)&i=1\\ \frac{1}{n}\sum_{j=1}^{n}\left(\left[\pi(x_{j},\theta)-(1-\pi(x_{j},\theta))\frac{z}{1-z}\right]\right)&i=2\end{array}\right.$ (5)
where $\pi(x_{j},\theta)$ and $\pi\left(x_{j},\hat{\theta}\right)$ are the true and predicted risk of the $j^{th}$ patient, respectively.
@@ 2.3 Expected value of current information
The true NBs formulated in the previous section (equation 4) all depend on the true parameters $\eta$ and $\theta$, and their estimate expressed in 5, after substituting $\eta$ by its empirical distribution, still depends on $\theta$. However, we do not know the true value of $\eta$ and $\theta$, and instead know about their likely values via $P(\eta,\theta|d)$. One can compute the expected NB (ENB) of the three strategies with respect to this distribution as
$ENB(i;d)=\mathbf{E}_{\eta,\theta|d}\left[NB\left(i;\hat{\theta},\eta,\theta\right)\right]$ (6)
for $i=0,1,2$, where $E_{\eta,\theta|d}$ represents the expectation with respect to the posterior distribution of $P(\eta,\theta|d)$.
Consequently, decision making under current info can be formulated as follows: given sample $d$, a decision maker fits a model to obtain the model parameter estimate $\hat{\theta}$, estimates the ENB of the model, and declares a winner $W(d)$, which returns a value in $\{0,1,2\}$. The most rational approach for a risk-neutral decision maker is to choose the winning strategy as
$W(d)=\operatorname*{arg\,max}_{i}\{ENB(i;d)\}$ (7)
To compute $ENB(i;d)$ for model development, we propose a generic bootstrap-based algorithm as follows. First take a set of weights for the development sample $d$ as the Bayesian bootstrap sample $d^{(t)}=\left(x^{(t)},y^{(t)}\right)$, a weighted version of the original development sample. Next, refit the same original risk prediction model (a weighted model) on $d^{(t)}$ to obtain the model coefficient estimates $\theta^{(t)}$ and consequently, the vector of predicted risks $\pi^{(t)}=\pi\left(x^{(t)},\theta^{(t)}\right)$ from the fitted model. We treat $d^{(t)}$ as the population and $\theta^{(t)}$ as the vector of true model parameters, and $\pi^{(t)}$ as the vector of true risks. One can plug in these values in (5) to obtain the NB of different strategies on the development sample. To estimate ENBs and consequently compute EVSI, we simply need to repeat the above steps $T$ times, $t=1,\dots,T$, and then use the average of NBs as their estimated expectation, that is
$\hat{ENB}(i;d)=\left\{\begin{array}[]{ll}0&i=0\text{ (treat none)}\\ \frac{1}{T}\sum_{t=1}^{T}\hat{NB}\left(1;\hat{\theta},\theta^{(t)}\right)&i=1\text{ (use model)}\\ \frac{1}{T}\sum_{t=1}^{T}\hat{NB}\left(2;\hat{\theta},\theta^{(t)}\right)&i=2\text{ (treat all)}\end{array}\right.,$ (8)

### v2

@@ Abstract
Risk prediction models are often advertised as deterministic functions that map covariates to predicted risks. However, they are typically trained using finite samples, and as such, their predictions are inherently uncertain. This uncertainty has been addressed in terms of uncertainty around metrics of model performance (e.g., confidence intervals around c-statistic), as well as uncertainty or instability of predictions. Correspondingly, sample size calculations for model development studies target the precision of estimates of summary statistics and the stability of predictions. However, when evaluating the clinical utility of a model (as in Net Benefit (NB) calculations in decision curve analysis), statistical inference is less relevant. From a decision-theoretic perspective, the finite size of the sample results in utility loss due to the discrepancy between the fitted model and the correct model. From this perspective, procuring more development data is associated with an expected gain in the utility of using the model. In this work, we define the Expected Value of Sample Information (EVSI) as the expected gain in clinical utility, defined in NB terms, by procuring an additional development sample of a given size. We propose a bootstrap-based algorithm for EVSI computations and demonstrate its feasibility and face validity in a case study. We conclude that decision-theoretic metrics can complement classical inferential methods when designing studies aimed at developing risk prediction models. predictive analytics; precision medicine; decision theory; value of information; bayesian statistics
@@ 1 Introduction
Risk prediction models that utilize patient characteristics to predict the risk of clinically relevant events are important enablers of precision medicine and individualized care (Kattan and others, 2016; Steyerberg, 2019; Van Calster and others, 2013). Typically, such models are classical regression-based or machine learning algorithms developed using a finite ‘development’ sample from a relevant target population. Examples of such models are the Framingham Risk Score for predicting the risk of cardiovascular outcomes (D’Agostino Sr and others, 2008), or the Adnex score for predicting the risk of malignancy in ovarian tumors (Timmerman and others, 2016).
@@
Predictions from such models are inherently subject to uncertainty, primarily due to the finite size of the development sample (Harrell, 2015; Steyerberg, 2019). Conventionally, such uncertainty is communicated through standard frequentist methods around performance metrics of the model (e.g., 95% confidence interval around c-statistic or calibration line (DeLong and others, 1988)). Recent developments have focused on studying the stability of predicted risks, with proposals on the graphical and statistical presentations of variation in predicted risks (Riley and Collins, 2023). Similarly, in designing risk prediction development studies, contemporary approaches for sample size calculations are based on statistical inference on classical metrics of model performance. For example, Riley and others (2024) have proposed multi-criteria sample size rules targeting pre-specified interval widths around prediction error, calibration, and discrimination. These developments have provided much needed objectivity in the design of risk prediction development and validation studies.
In addition to classical metrics of model performance around calibration and discrimination, evaluating the clinical utility of a model, expressed in terms of net benefit (NB), has become a standard component of model performance (Vickers and Elkin, 2006a; Efthimiou and others, 2024), with recent best practice recommendations emphasizing the calculation of NB at risk threshold(s) of interest when developing a new model (Efthimiou and others, 2024).
@@
Among conventional metrics of model performance, NB is a decision-theoretic one, in the sense that the use of a risk prediction model is clinically justified if its expected NB is above the expected NB of the alternative decisions. Decision theory provides a fundamentally different approach towards uncertainty quantification than methods based on conventional statistical inference (Claxton, 1999). From the perspective of this theory, uncertainty results in the discrepancy between the predicted and correct (calibrated) risks, which in turn might lead to an incorrect medical decision (a decision based on predicted risk that is different from when knowing the correct risk). Thus, from this perspective, uncertainty can be associated with a loss of clinical utility (Felli and Hazen, 1998). Such a decision-theoretic approach towards evaluating the consequences of uncertainty in decisions, generally referred to as Value of Information (VoI) analysis, has been employed in various fields, including decision-analytic modeling for the economic evaluation of health technologies (Jackson and others, 2022, 2019), environmental risk analysis (Yokota and Thompson, 2004; Weinstein, 1980; O’Hagan and others, 2006), and data-driven cost-effectiveness analyses of clinical trials (Everest and others, 2023). Recently, VoI methods have been adapted to NB calculations for risk prediction models. Specifically, the expected value of perfect information (EVPI) has been applied to both development and validation phases of risk prediction models (Sadatsafavi and others, 2022, 2023). In the context of model development, EVPI quantifies the expected utility gain over a model developed using a finite sample by completely knowing the correct (strongly calibrated) model.
@@
However, uncertainty can never be entirely eliminated. Procuring more samples for model development is expected to reduce, but not eliminate, prediction uncertainty. As such, the EVPI puts an upper limit on the expected gain from future development studies. In decision theory, the expected value of sample information (EVSI) quantifies the potential value of collecting additional data to inform a specific research design (Claxton, 1999; Willan and Briggs, 2006). In health policy-making, EVSI has been applied to quantify the expected utility (net benefit) gain from conducting a clinical trial of a given design and sample size (Ades and others, 2004).
@@
The EVSI has recently been proposed for the validation phase of a model, where it quantifies the expected NB gain by reducing uncertainty in our estimation of the NB of a pre-specified model in a new population (Sadatsafavi and others, 2025). To the best of our knowledge, it has not yet been applied to the development phase of such models, where the sample is used for model fitting and NB loss occurs because of the discrepancy between the fitted model and the correct model that maps predictors to the correct conditional probability of outcome risk. A particular utility of EVSI calculations in this context is at the end of the development phase of a model, where a decision needs to be made whether procuring more development data is required, or the model should move on to the next phase (e.g., validation or impact analysis). In addition to the development phase of a model, in many instances when a model is validated in a new population, some form of model updating will be required ranging from changing the intercept, to refitting the model in the new population (Binuya and others, 2022). This means in many practical instances, the investigators need to assess the NB of their model in the same data that is used to fit or update the model. As such, EVSI calculations for model development can have wide applicability.
@@ 2 Context and Notations
We focus on the development phase of a risk prediction model for a binary outcome (e.g., whether an individual with a recent heart attack will die in the next 30 days). We assume the investigator has had access to $n$ independent and identically distributed observations from the relevant target population to develop a prediction model for the conditional risk of the outcome given a set of predictor values. Due to the finite size of the development sample, predictions from such a model might not be accurate even if the model form is correctly specified. Procuring a further development sample can improve the accuracy of predictions, and reduce such a loss. The figure of merit is the expected gain in clinical utility from procuring a future development sample $D^{*}$ of size $n^{*}$ from the same target population. We assume the future sample will be exchangeable with the current sample (thus the investigator will naturally combine the current and future samples to refit their model). We quantify clinical utility in terms of NB, as quantified in decision curve analysis (DCA - Vickers and Elkin 2006b). An overview of NB calculations are provided below.
@@ 2.1 Notations
Let $\mathbf{X}$ represent the vector of predictors and $Y$ the binary outcome, and $d=(\mathbf{x},y)=\{(\mathbf{x_{i}},y_{i})\}_{i=1}^{n}$ be the already procured sample that is used to develop the existing model. The development data are presumed to obey the following joint distribution
$P_{\theta,\eta}(\mathbf{X},Y)=P_{\eta}(\mathbf{X})P_{\theta}(Y|\mathbf{X}),$ (1)
where $\eta$ and $\theta$ are the parameters indexing, respectively, the marginal distribution of $\mathbf{X}$ (i.e., the case-mix) and the conditional distribution of $Y$ given $\mathbf{X}$ ($Y|\mathbf{X}$, i.e., the response model). Let $\pi(\mathbf{x};\theta)=P(Y=1|\mathbf{X}=\mathbf{x})$ be the true risk of outcome for patient with characteristics $\mathbf{x}$, and consequently, $\pi(\mathbf{X};\theta)=P(Y=1|\mathbf{X})$ be a random draw from the distribution of true risks specified by the true model with parameter vector of $\theta$. Related, $\hat{\eta}$ and $\hat{\theta}$ represent the estimated case mix and coefficients of the prediction model, and therefore, $\pi(\mathbf{x};\hat{\theta})$ returns the predicted risk of outcome for patient with characteristics $\mathbf{x}$ using the fitted model. In what follows, we parametrically model $Y$, estimating $\theta$ (regression coefficients) in a general regression framework, while we adopt a non-parametric approach towards modeling case-mix, taking $\hat{\eta}$ as the empirical distribution of $\mathbf{X}$.
@@ 2.2 Net benefit
We determine clinical utility by using NB, as performed in DCA. Details of NB calculations are provided elsewhere (Vickers and Elkin, 2006a). In brief, when converting continuous risk predictions into binary treatment decisions, it is necessary to set a treatment threshold on predicted risks. This threshold should ideally reflect the relative importance of avoiding unnecessary treatments (false positives) versus ensuring necessary treatments (true positives).
Among the proportion of the population whose predicted risk is above the threshold, a proportion will experience the event, constituting the True Positive cases. The rest of this population will not experience the event, thus comprising the False Positive population. Compared to the strategy of not treating anyone (with a default NB of zero), the use of the model confers utility gain by identifying True Positives who will appropriately receive the treatment, and causes utility loss through False Positives who will unnecessarily receive the treatment. As such, the NB of using the model can be computed as:
$\text{NB}=P(\text{True Positive})-wP(\text{False Positive}),$ (2)
where $w$ represents the exchange rate between the utility gain of True Positives and utility loss of False Positives.
Vickers and Elkin (2006a) showed that the value of $w$ can be inferred from the risk threshold itself. Specifically, to convert a continuous predicted risk into a binary classification for treatment decisions, a decision-maker must set a risk threshold $z$. Individuals with predicted risks above $z$ are classified as high-risk and receive the treatment. Consider the cohort of individuals with outcome risk precisely equal to $z$. Among these individuals, the decision-maker is ambivalent between the treatment and no treatment decisions, implying that the utility of treating this cohort is zero. Because among this cohort, $P(TruePositive)=z$ and $P(FalsePositive)=1-z$, setting the utility to zero shows $w=\frac{z}{1-z}$. In practice, the NB is often calculated for a plausible range of thresholds (Sadatsafavi and others, 2025).
To make a decision whether the model is useful in this population, we shall compare its NB against the NB of at least two default decisions: treating all and treating none. The latter has 0 NB by definition, and the NB of treating all can be written as (noting that treating all is akin to using a model that classifies everyone as high risk)
@@
[…]
$NB\left(i;\hat{\theta},\eta,\theta\right)=\left\{\begin{array}[]{ll}0&i=0\text{ (treat none)}\\ \mathbb{E}\left\{I\left(\pi\left(\mathbf{X},\hat{\theta}\right)\geq z\right)\left[\pi(\mathbf{X},\theta)-(1-\pi(\mathbf{X},\theta))\frac{z}{1-z}\right]\right\}&i=1\text{ (use model)}\\ \mathbb{E}\{\pi(\mathbf{X},\theta)-(1-\pi(\mathbf{X},\theta))\frac{z}{1-z}\}&i=2\text{ (treat all)}\end{array}\right.,$ (4)
@@
Assume for the moment that we have access to the true parameters $\theta$. While plugging in the empirical estimate of $\eta$ (a weight of $1/n$ assigned to each observation), then a consistent estimate of NBs based on the development sample $d$ is:
$\hat{NB}\left(i;\hat{\theta},\theta\right)=\left\{\begin{array}[]{ll}0&i=0\\ \frac{1}{n}\sum_{j=1}^{n}\left(I\left(\pi\left(\mathbf{x}_{j},\hat{\theta}\right)>z\right)\left[\pi(\mathbf{x}_{j},\theta)-(1-\pi(\mathbf{x}_{j},\theta))\frac{z}{1-z}\right]\right)&i=1\\ \frac{1}{n}\sum_{j=1}^{n}\left(\left[\pi(\mathbf{x}_{j},\theta)-(1-\pi(\mathbf{x}_{j},\theta))\frac{z}{1-z}\right]\right)&i=2\end{array}\right.$ (5)
where $\pi(\mathbf{x}_{j},\theta)$ and $\pi\left(x_{j},\hat{\theta}\right)$ are the true and predicted risk of the $j^{th}$ patient, respectively. The above NB estimates depend on the true value of $\theta$, which is unknown. In the next section, we will propose a Bayesian bootstrap-based approach to address this issue. Note that the ordinary bootstrap can be interpreted as a (approximate nonparametric) Bayesian bootstrap procedure, where the empirical distribution function of the data plays a role analogous to a non-informative prior, and the resampling process approximates the posterior distribution.
@@ 2.3 Expected value of current information
The true NBs formulated in the previous section (Equation 4) all depend on the true parameters $\eta$ and $\theta$. We do not know the true value of the parameters, and instead know about their likely values via their posterior distribution. One can compute the expected NB (ENB) of the three strategies with respect to this distribution as
$ENB(i;d)=\mathbf{E}_{\theta,\eta|d}\left[\hat{NB}\left(i;\hat{\theta},\theta,\eta\right)\right]$ (6)
for $i=0,1,2$, where $\mathbf{E}_{\theta,\eta|d}$ represents the posterior expectation of (4) with respect to the joint posterior distribution on $(\theta,\eta)$. By using the Bayesian bootstrap method to estimate the above posterior expectation (detailed in the next section), we assumed posterior independence between $\theta$ and $\eta$. This property would hold immediately if both parameters were treated parametrically and prior independence were assumed. Consequently, (6) can be simplified by omitting $\eta$, yielding $ENB(i;d)=\mathbf{E}_{\theta|d}\left[\hat{NB}\left(i;\hat{\theta},\theta\right)\right]$, where $\hat{NB}\left(i;\hat{\theta},\theta\right)$ is defined in (5).
Consequently, decision making under current information can be formulated as follows: given sample $d$, a decision maker fits a model to obtain the model parameter estimate $\hat{\theta}$, estimates the ENB of the model as well as other baseline strategies (as explained below), and declares a winner $W(d)$, which returns a value in $\{0,1,2\}$
$W(d)=\operatorname*{arg\,max}_{i}\{ENB(i;d)\},$ (7)
which leads us to the ENB of the current information being $ENB\left(W(d);d\right)$.
To compute $ENB(i;d)$ for model development, we propose a bootstrap-based algorithm as follows. First take a set of weights for the development sample $d$ as the Bayesian bootstrap sample $d^{(t)}=\left(\mathbf{x}^{(t)},y^{(t)}\right)$, a weighted version of the original development sample. Next, we refit the same original risk prediction model on $d^{(t)}$ to obtain the model coefficient estimates $\theta^{(t)}$ and consequently, predicted risks $\pi^{(t)}=\pi\left(\mathbf{x}^{(t)},\theta^{(t)}\right)$ from the fitted model. We treat $d^{(t)}$ as the population, $\theta^{(t)}$ as the vector of true model parameters, and $\pi^{(t)}$ as the vector of true risks. One can plug in these values in (5) to obtain the NB of different strategies on the original development sample. To estimate ENBs and consequently compute EVSI, we simply need to repeat the above steps $T$ times, $t=1,\dots,T$, and then use the average of NBs as their estimated expectation, that is
$\hat{ENB}(i;d)=\left\{\begin{array}[]{ll}0&i=0\text{ (treat none)}\\ \frac{1}{T}\sum_{t=1}^{T}\hat{NB}\left(1;\hat{\theta},\theta^{(t)}\right)&i=1\text{ (use model)}\\ \frac{1}{T}\sum_{t=1}^{T}\hat{NB}\left(2;\hat{\theta},\theta^{(t)}\right)&i=2\text{ (treat all)}\end{array}\right..$ (8)
The bootstrap algorithm we propose for estimating the above ENBs mirrors the general bootstrap-based approach for computing optimism-corrected NB described by Harrell (2015). Because of this shared bootstrap foundation, our method can be readily embedded within the optimism-adjusted, or drawing instability plots for model development (Riley and Collins, 2023).